OpenStudy (anonymous):

x-3=3/11 can you show me the steps?

OpenStudy (anonymous):

Just add \(3\) to both sides and then simplify.

OpenStudy (anonymous):

x-3=3/11 +3| +3 x=3/11+3 x=3 and 3/11

OpenStudy (anonymous):

turn 3 into 11ths. if 1 in 11ths is 11/11, then 3 is 2 times that, or 33/11, so that is the same as x-33/11=3/11 then add 33/11 and you get x=36/11 or x=3 and 3/11

OpenStudy (anonymous):

haha and sorry i meant that 3 is THREE times that, or 33/11, not 2 times that. sorry